# Large Language Model Unlearning via Embedding-Corrupted Prompts

### v1

@@ C.1.4 Why Not Membership Inference Attacks (MIAs)?
In this paper, we follow most prior work on LLM unlearning, which generally do not use membership inference attack (MIA) methods to evaluate the effectiveness of unlearning for LLMs [57, 65, 53, 22, 34, 87, 136, 135, 80, 70, 142, 54, 59].
We do not consider MIA methods to evaluate our models for three major reasons. First, state-of-the-art MIAs require training multiple (up to hundreds) shadow models [21] on subsets of the entire training set, which are not feasible for LLMs’ setting, which requires access to the pre-training data or fine-tuning a large number of models on subsets of of the fine-tuning data. MIAs without training shadow models have been demonstrated to overestimate the effectiveness of unlearning [51] due to the non-uniform difficulty of learning/unlearning each sample.
Second, evidence suggests that existing MIAs for LLMs, even the state-of-the-art one [116], generally barely perform better than random guessing due to both training on large pre-training data size for a small number of iterations and the fuzzy boundary between members and non-members [32].
Third, as stated in Section 2.2, we do not consider the privacy aspect of unlearning in this work, and our threat model does not include privacy risks. Knowing if a single sample is a membership also does not significantly increase the risk in our threat model.
Additionally, performing such MIAs typically require at least the model internal states [32] (e.g., activations), which is not in the scope of our threat model (i.e, only text output and logits).
In fact, the forget quality metric described in Section C.1.1 and the probing evaluation in Section C.1.2 aligns with the goal of MIAs. The forget quality assesses if the forget set distributions on the unlearning model and the retained models can be distinguished. The linear probe tries to infer the correct answers from the model output, assuming that the the accuracy of the linear probe on a retained model is random-guessing level. Obtaining the same accuracy might imply indistinguishability.
@@ C.2 Preparing LLMs for Unlearning
We use the original code [^7] https://github.com/locuslab/tofu provided alongside the TOFU dataset [87] for fine-tuning to ensure consistency. Following their experimental setup, we fine-tune two models, Phi-1.5 [71] and Llama-2-7B-Chat [124], on the entire TOFU dataset to obtain the model to be subjected to unlearning. Following the retain/forget splits provided in the dataset, we fine-tune each model on each of the three different splits, 99%, 95%, and 90% of the full dataset, excluding the forget data, to obtain the retained models. These three splits also correspond to unlearning 1%, 5%, and 10% of the samples, respectively. We employ the same hyperparameters as provided in both the paper and the accompanying code. Both models are trained with a batch size of 4, accumulating gradients for 4 steps on 2 NVIDIA A6000 GPUs, resulting in an effective batch size of 32, with a learning rate of 1e-5 for Llama-2-7B-Chat and 2e-5 for Phi-1.5. For the negative preference optimization [142] baselines, we follow a similar procedure and uses the code provided by the original authors [^8] https://github.com/licong-lin/negative-preference-optimization.
@@
The knowledge assessment of all multiple-choice questions in WMDP [70] and MMLU [52] subsets is performed directly on the pre-trained models (or models unlearned from the pre-trained checkpoints for unlearning evaluation). Therefore, we do not fine-tune models based on the multiple-choice questions for the the WMDP unlearning task.
@@
For the copyrighted content unlearning task, we first verify that all the considered LLMs cannot generate the original corpus. For HP Book, while some parts of the text corpus could potentially be included during pretraining, we see little sign of generating the text spans verbatim for all models we considered in the copyrighted content unlearning experiments. This is also reflect in the low similarity scores from tables in Section E.5. For BBC News articles, we only consider articles published in February 2024, which is beyond the knowledge cutoff of most models considered. We fine-tune them on the copyrighted content corpus to ensure that they are able to generated the original passage. For the HP book, we split the text into chunks of up to 256 tokens (based on the tokenization scheme used for the specific model). For BBC News articles, we concatenate the news title with the news content with a single space in the middle. The title of the news article is used as the prompt for generation. To ensure that our models can indeed generate the copyrighted content, we fine-tune all models on the two text corpora for 5 epochs, using a batch size of 4 and a learning rate of 2e-5 on two NVIDIA A100 GPUs.
@@ C.3 Prompt Classifiers
Table 5: The false negative rate (FNR) and false positive rate (FPR) of the prompt classifiers without thresholding.
@@
WMDP (All) 0.0 0.0 0.0 0.0 0.0
$\text{WMDP}_{\text{Synthetic}}$ (All) 0.0 0.0 0.0 0.0 0.0004
$\text{WMDP}_{\text{o.o.d}}$ (All) 0.0 0.0 0.2721 0.016 0.003
MMLU (Economics) 0.0 0.0 0.0 0.0 0.0
MMLU (Physics) 0.0 0.0 0.0 0.0 0.0
MMLU (Law) 0.0 0.0 0.0 0.0 0.0004
@@
WMDP (All) 0.0 0.0 0.0 0.0006 0.0004
$\text{WMDP}_{\text{Synthetic}}$ (All) 0.0 0.0 0.0 0.0006 0.0047
$\text{WMDP}_{\text{o.o.d}}$ (All) 0.0 0.0 0.1267 0.1016 0.0834
MMLU (Economics) 0.0 0.0 0.0 0.0 0.002
MMLU (Physics) 0.0 0.0 0.0 0.0 0.001
MMLU (Law) 0.0 0.0 0.0 0.0 0.001
@@ C.3.1 Prompt Classifiers’ Training Datasets
We train a single classifier to classify multiple choice questions in all three subjects of the WMDP datasets (i.e., biology, chemistry, and cybersecurity). We format the question strictly following the original evaluation [70] using the Language Model Evaluation Harness [40] style in LABEL:lst:wmdp_template, where the {subject}, {question}, and {choice_#} fields are replaced by the actual text for each multiple-choice question. This template is also used as the prompt template for the main LLM to make predictions. For negative samples, we use a combination of the auxiliary training set and the development set of MMLU [52]. We do so because the questions in the auxiliary training set do not have subjects, which might result in a shortcut picked up by the classifier. We use the sampels in development set (with subjects) to mitigate that shortcut. Note that MMLU has its own validation set, so we believe that using the development set (with only 285 samples) is a reasonable choice.
@@
Different from TOFU and two other copyrighted content dataset, we require the classifier trained on the WMDP questions to generalize to unseen questions in relevant domains, while not flagging relevant questions in similar domains (e.g., virology, high school chemistry, and computer security in MMLU) as positive. Additionally, the method, RMU, proposed along with the WMDP dataset does not require access to the forget questions. To avoid fully relying on the WMDP questions to train the classifier, we first restrict ourselves to access only 10% of the WMDP questions, selected randomly before training and development. We also subsample a fixed set of 3K samples (out of 99.8K) in the auxiliary training set of MMLU to train and test the false positive rate of the prompt classifier. Note that, the RMU method proposed in [70] does not require access to any of the questions. Therefore, we also incorporate a setting where we train a prompt classifier on only synthetic data. The synthetic dataset has only 300 questions (100 for each of the biology, chemistry, and cybersecurity subsets), and all multiple-choice questions are generated by GPT-4 [2] to resemble the style and difficulty of real WMDP questions. In this way, we align with RMU’s setup, which does not access to real questions during model development. For a detailed split, please see Table 4. The perfomrnace of both prompt classifiers are reported in Table 6. We see that the WMDP prompt classifier trained on only synthetic data have almost the same performance as one trained on 10% of the real questions.
@@
In Tables 5, 6 and 7, we also show an o.o.d. setup, where we assume we only have 100 biology questions, 100 cybersecurity questions, but only 2 chemistry questions. In this scenario, the false negative rates of the original classifier and the simple-threshold classifier are high due to the lack of training samples for chemistry. In Table 7, we demonstrate that conformal prediction can alleviate this issue via including more uncertain samples as negative, which reduces the false negative rate by approximately 14%.
@@
We train a separate classifier for each of the unlearning category: economics, physics, and law. This is because for each task, we hope to unlearn only the selected category and retain the rest of the categories as defined in [70], which makes training a single prompt classifier infeasible experiment-wise. Similar to the WMDP setting, we only restrict ourselves to have access to the development set of MMLU, which only contains 10-15 samples for the forget target. As shown in Table 6, all classifiers achieve perfect predictions, except that our MMLU law classifier has 10 false positives: 2 in HellaSwag and 8 in BoolQ. After manual inspection of the false positive samples, we observed that all of them are directly related to law and closely resemble the style of questions in the MMLU’s law category. We believe this is due to the generalization capacity of the classifier, which is difficult to avoid given the high similarity in those questions. In fact, after giving the sample questions to humans, they cannot tell the difference between the MMLU law questions and HellaSwag/BoolQ law questions. Therefore, seeing that the false positive rate is still reasonably low (i.e., 0.0004), we consider this type of error to be tolerable.
@@
Since the goal is to prevent users from obtaining the copyrighted content by training data extraction, we purchased the Harry Potter and the Sorcerer’s Stone [106] ebook and extracted the entire corpus to train our HP Book prompt classifier. We split the entire book into sentences using spaCy sentencizer [^9] https://spacy.io/api/sentencizer, and only select sentences with more than ten characters. The sentences with ten or fewer characters are mostly neutral sentences, line escape, whitespaces, or other punctuation. All remaining sentences are considered as positive sentences for our prompt classifier. For the negative samples, we employ the BookMIA dataset [^10] https://huggingface.co/datasets/swj0419/BookMIA [116], which contains more than 9K text snippets from a wide range of real books. The BookMIA dataset contains some snipepts from Harry Potter and the Sorcerer’s Stone [106], which are all removed before training. In this task, we do not require our prompt classifier to generalize, so we do not use a test forget set. We subsample and split the retain set (i.e., BookMIA) into two equal-sized sets and use them to train/test the performance of the prompt classifier. In Table 6, while our classifier incorrectly predicts several samples, by manual examination, we find that those samples are mostly noise (i.e., neutral sentences).
@@
We use BBC News articles [^11] https://huggingface.co/datasets/RealTimeData/bbc_news_alltime published in February 2024 as the positive samples and 9K+ news articles in the English language from the CC-News dataset [49] as the negative samples to train the prompt classifier. To prevent shortcut being learned, we format both datasets in the same way and remove the “ - BBC ###” suffix in the title of the BBC News dataset. The prompt classifier is only trained on the title of the news article. To prevent more sophisticated extraction attacks, one can train the prompt classifier using news articles split into sentences, similar to the HP Book dataset. However, due to the long lengths of the full news articles, we only consider classification of the news title.
@@
Most prior work only evaluate the retain ability of the unlearned LLM using the retain set associated with the unlearning task. However, the results reported on the retain set might not fully reflect the general utility in the real-world setting. This is because the retain set, while being disjoint with the forget set, might still have a similar distribution to the forget set in some aspect. Therefore, instead of solely relying on the regular retain set, we consider a large set of out-of-distribution samples to measure general utility. In the general set, we eleven common LLM benchmarks listed in Table 8: MMLU [52], ARC-Easy [26], ARC-Challenge [26], OpenBookQA [89], HellaSwag [139], Winogrande [108], TruthfulQA [75], CommonsenseQA [120], PIQA [15], SocialIQA [111], and BoolQ [25]. They amount to 41,297 samples in total. We evaluate all prompt classifiers (from all datasets mentioned above) on the general set after tuning the parameters of the prompt classifiers. For all datasets, we use the test set if the labels are publicly available; otherwise, we use the validation set. For TruthfulQA, we use the MC1 subset for evaluation.
@@ C.3.2 Training A prompt classifier
For all prompt classifiers used for prompt content detection, we choose RoBERTa-base [79] as the base model for fine-tuning. The hyperparmaeters are selected following a prior work that improves stability during training [92]. Since in most cases the number of positive samples (the forget samples) is much less than the negative samples, we reweight the class-wise losses using the inverse frequency. Once the optimal number of epochs is determined, we fine-tune the model again on the combined training and validation set and use it as the final prompt classifier used in inference-time unlearning. We also evaluate all prompt classifiers on the general set outlined in Table 8.
@@
We report the performance of the original prompt classifier, classifier with simple thresholding, and classifier with conformal prediction in Tables 5, 6 and 7. We demonstrate that in all settings considered, our best prompt classifiers have achieved satisfying performance reflected by the low false negative rate on the forget set and low false positive rate on the retain set. On the general set (i.e., the suite of NLP benchmarks), most prompt classifiers have zero false positive predictions, suggesting that the performance on the main LLM on samples irrelevant to the forget set is unlikely to be affected.

### v2

@@ C.1.4 Why Not Membership Inference Attacks (MIAs)?
In this paper, we follow most prior work on LLM unlearning, which generally does not use membership inference attack (MIA) methods to evaluate the effectiveness of unlearning for LLMs [60, 68, 56, 23, 36, 90, 143, 142, 83, 73, 149, 57, 62].
We do not consider MIA methods to evaluate our models for three major reasons. First, state-of-the-art MIAs require training multiple (up to hundreds) shadow models [22] on subsets of the entire training set, which is not feasible in the LLM setting, as it requires access to the pre-training data or fine-tuning a large number of models on subsets of the fine-tuning data. MIAs without training shadow models have been demonstrated to overestimate the effectiveness of unlearning [54] due to the non-uniform difficulty of learning/unlearning each sample.
Second, evidence suggests that existing MIAs for LLMs, even the state-of-the-art ones [122], generally barely perform better than random guessing due to both training on large pre-training datasets for a small number of iterations and the fuzzy boundary between members and non-members [33]. In addition, recent work [91] shows that Min-KProb leads to: 1) significant variance based on the random selection of datasets used for evaluation, 2) improved results when the two subsets (in this case, the forget and holdout sets) are drawn from different distributions, and 3) an empirical overestimation of false positives. This last point indicates that the distribution gap (such as a temporal shift, also noted by [33]) introduces a confounding variable in the discrimination process, as the forget and holdout sets might vary in more than one aspect.
Third, as stated in Section 2.2, we do not consider the privacy aspect of unlearning in this work, and our threat model does not include privacy risks. Knowing whether a single sample is a member also does not significantly increase the risk in our threat model.
Additionally, performing such MIAs typically requires at least the model’s internal states [33] (e.g., activations), which are not within the scope of our threat model (i.e., only text output and logits).
In fact, the forget quality metric described in Section C.1.1 and the probing evaluation in Section C.1.2 align with the goal of MIAs. The forget quality assesses whether the forget set distributions on the unlearned model and the retained models can be distinguished. The linear probe tries to infer the correct answers from the model output, assuming that the accuracy of the linear probe on a retained model is at the random-guessing level. Achieving the same accuracy might imply indistinguishability.
@@ C.2 Preparing LLMs for Unlearning
We use the original code [^7] https://github.com/locuslab/tofu provided alongside the TOFU dataset [90] for fine-tuning to ensure consistency. Following their experimental setup, we fine-tune two models, Phi-1.5 [74] and Llama-2-7B-Chat [131], on the entire TOFU dataset to obtain the model to be subjected to unlearning. Following the retain/forget splits provided in the dataset, we fine-tune each model on each of the three different splits—99%, 95%, and 90% of the full dataset, excluding the forget data, to obtain the retained models. These three splits also correspond to unlearning 1%, 5%, and 10% of the samples, respectively. We employ the same hyperparameters as provided in both the paper and the accompanying code. Both models are trained with a batch size of 4, accumulating gradients for 4 steps on 2 NVIDIA A6000 GPUs, resulting in an effective batch size of 32, with a learning rate of 1e-5 for Llama-2-7B-Chat and 2e-5 for Phi-1.5. For the negative preference optimization [149] baselines, we follow a similar procedure and use the code provided by the original authors [^8] https://github.com/licong-lin/negative-preference-optimization.
@@
The knowledge assessment of all multiple-choice questions in WMDP [73] and MMLU [55] subsets is performed directly on the pre-trained models (or models unlearned from the pre-trained checkpoints for unlearning evaluation). Therefore, we do not fine-tune models based on the multiple-choice questions for the WMDP unlearning task.
@@
For the copyrighted content unlearning task, we first verify that all the considered LLMs cannot generate the original corpus. For the HP Book, while some parts of the text corpus could potentially be included during pretraining, we see little sign of generating the text spans verbatim for all models we considered in the copyrighted content unlearning experiments. This is also reflected in the low similarity scores from tables in Section E.5. For BBC News articles, we only consider articles published in February 2024, which is beyond the knowledge cutoff of most models considered. We fine-tune them on the copyrighted content corpus to ensure that they are able to generate the original passage. For the HP book, we split the text into chunks of up to 256 tokens (based on the tokenization scheme used for the specific model). For BBC News articles, we concatenate the news title with the news content, with a single space in the middle. The title of the news article is used as the prompt for generation. To ensure that our models can indeed generate the copyrighted content, we fine-tune all models on the two text corpora for 5 epochs, using a batch size of 4 and a learning rate of 2e-5 on two NVIDIA A100 GPUs.
@@ C.3 Prompt Classifiers
Table 5: The false negative rate (FNR) and false positive rate (FPR) of the prompt classifiers without thresholding. If the FNR of $D^{\textrm{Test}}_{f}$ is not reported, it means that the corresponding unlearning target does not require generalization outside the scope of the forget set. The $D_{g}$ set contains out-of-distribution prompts from eleven NLP benchmarks listed in Table 9.
@@
WMDP (All) 0.0021 0.0056 0.04 0.004 0.01
$\text{WMDP}_{\text{Synthetic}}$ (All) 0.0048 0.009 0.009 0.0066 0.007
$\text{WMDP}_{\text{o.o.d}}$ (All) 0.0009 0.0004 0.2721 0.016 0.003
MMLU (Economics) 0.0002 0.0008 0.0006 0.0024 0.0219
MMLU (Physics) 0.0005 0.0003 0.0143 0.003 0.072
MMLU (Law) 0.0007 0.0001 0.0048 0.017 0.087
@@
WMDP (All) 0.0003 0.0007 0.0002 0.018 0.0005
$\text{WMDP}_{\text{Synthetic}}$ (All) 0.0005 0.0006 0.0001 0.007 0.049
$\text{WMDP}_{\text{o.o.d}}$ (All) 0.0004 0.0002 0.1253 0.1021 0.0839
MMLU (Economics) 0.0021 0.00556 0.0046 0.0039 0.023
MMLU (Physics) 0.001 0.0034 0.0062 0.0036 0.012
MMLU (Law) 0.0035 0.0026 0.0071 0.051 0.014
@@
Category Subcategory (Retain) False Positive
Economics Business Ethics 0/100
Econometrics 0/114
Management 0/103
Marketing 0/234
Professional Accounting 2/282
Law Business Ethics 0/100
Jurisprudence 8/108
Logical Fallacies 0/163
US Foreign Policy 0/100
Physics Abstract Algebra 0/100
College Mathematics 0/100
Electrical Engineering 2/145
Elementary Mathematics 0/378
Formal Logic 0/126
High School Mathematics 2/270
High School Statistics 0/216
Category Subcategory (Forget) False Negative
Economics High School Microeconomics 0/238
High School Macroeconomics 0/390
Law Professional Law 8/1534
International Law 2/121
Physics High School Physics 3/151
Conceptual Physics 4/235
College Physics 0/102
Table 8: False positive and false negative of the MMLU classifiers on highly-related subjects.
@@ C.3.1 Prompt Classifiers’ Training Datasets
We train a single classifier to classify multiple-choice questions in all three subjects of the WMDP dataset (i.e., biology, chemistry, and cybersecurity). We format the question strictly following the original evaluation [73] using the Language Model Evaluation Harness [43] style in LABEL:lst:wmdp_template, where the {subject}, {question}, and {choice_#} fields are replaced by the actual text for each multiple-choice question. This template is also used as the prompt template for the main LLM to make predictions. For negative samples, we use a combination of the auxiliary training set and the development set of MMLU [55]. We do so because the questions in the auxiliary training set do not have subjects, which might result in a shortcut learned by the classifier. We use the samples in the development set (with subjects) to mitigate that shortcut. Note that MMLU has its own validation set, so using the development set (with only 285 samples) is a reasonable choice.
@@
Unlike TOFU and other copyrighted content datasets, we require the classifier trained on WMDP questions to generalize to unseen questions in relevant domains while not flagging relevant questions in similar domains (e.g., virology, high school chemistry, and computer security in MMLU) as positive. To avoid fully relying on the WMDP questions to train the classifier, we restrict ourselves to access only 10% of the WMDP questions, selected randomly before training and development. We also subsample a fixed set of 3K samples (out of 99.8K) from the auxiliary training set of MMLU to train and test the false positive rate of the prompt classifier. Note that the RMU method proposed in [73] does not require access to any of the questions.
We incorporate a setting where we train a prompt classifier on only synthetic data. The synthetic dataset contains 300 questions (100 for each of biology, chemistry, and cybersecurity), all generated by GPT-4 [2] to resemble the style and difficulty of real WMDP questions. This aligns with RMU’s setup, which does not access real questions during model development. For a detailed split, please see Table 4. The performance of both prompt classifiers is reported in Table 6. We observe that the WMDP prompt classifier trained on synthetic data performs almost identically to one trained on 10% of the real questions. However, the authors of [129] identified that our initial classifiers overfit to the topic line. Specifically, since the topic line (LABEL:lst:wmdp_template) The following are multiple-choice questions (with answers) about {subject} appeared in the training data, the classifiers learned to classify forget and retain classes solely based on {subject}, rather than the question content. To address this, we used the official WMDP training corpora [73] to generate 3,000 synthetic problems in a style similar to the actual problems, using Llama-3.1-405B-Instruct [34]. We also included synthetic retain problems using the development set of MMLU for data generation. Including such data significantly reduced the false positive rate. We then fine-tuned a Llama-Guard-3-1B [34, 59] as the classifier, excluding the topic line during training.
@@
In Tables 5, 6 and 7, we also show an out-of-distribution (o.o.d.) setup where we assume access to only 100 biology questions, 100 cybersecurity questions, and 2 chemistry questions. In this scenario, the false negative rates of both the original and simple-threshold classifiers are high due to insufficient training samples for chemistry. In Table 7, we demonstrate that conformal prediction alleviates this issue by including more uncertain samples as negative, reducing the false negative rate by approximately 14%.
@@
We train a separate classifier for each unlearning category: economics, physics, and law. This is because, for each task, we aim to unlearn only the selected category while retaining the rest of the categories as defined in [73], making a single prompt classifier infeasible. Similar to the WMDP setup, we restrict ourselves to the MMLU development set, which contains only 10–15 samples for the forget target. We follow the same procedure used in the WMDP dataset and use the development set to generate synthetic forget and retain data to train the classifiers. To avoid overfitting, we also removed the topic line from all training samples. We report the detailed false positive and false negative statistics on the forget subject and related retain subjects in table 8.
@@
Since the goal is to prevent users from extracting copyrighted content through training, we purchased the Harry Potter and the Sorcerer’s Stone [112] ebook and extracted the corpus to train our HP Book prompt classifier. We split the book into sentences using spaCy’s sentencizer [^9] https://spacy.io/api/sentencizer, selecting only sentences with more than ten characters. Sentences with ten or fewer characters are mostly neutral sentences, line breaks, whitespace, or punctuation. The remaining sentences are treated as positive samples. For negative samples, we use the BookMIA dataset [^10] https://huggingface.co/datasets/swj0419/BookMIA [122], which contains over 9K text snippets from various real books. Snippets from Harry Potter and the Sorcerer’s Stone in the BookMIA dataset are removed before training. As generalization is not required, we do not use a test forget set. We split the retain set (i.e., BookMIA) into two equal-sized sets for training and testing the classifier’s performance. In Table 6, we find that the classifier misclassifies some samples, but manual examination reveals they are mostly neutral sentences.
@@
We use BBC News articles [^11] https://huggingface.co/datasets/RealTimeData/bbc_news_alltime published in February 2024 as positive samples and over 9K news articles in English from the CC-News dataset [52] as negative samples. To prevent shortcuts, we format both datasets consistently and remove the “ - BBC ###” suffix from the titles in the BBC News dataset. The prompt classifier is trained only on the titles of the news articles. To mitigate sophisticated extraction attacks, one could train the classifier using sentence-level splits, similar to the HP Book dataset. However, due to the length of full news articles, we focus solely on title-based classification.
@@
Most prior work only evaluates the retain ability of the unlearned LLM using the retain set associated with the unlearning task. However, the results reported on the retain set might not fully reflect general utility in real-world settings. This is because the retain set, while being disjoint from the forget set, might still share a similar distribution with the forget set in some aspects. Therefore, instead of solely relying on the regular retain set, we consider a large set of out-of-distribution samples to measure general utility. In the general set, we include eleven common LLM benchmarks listed in Table 9: MMLU [55], ARC-Easy [27], ARC-Challenge [27], OpenBookQA [94], HellaSwag [146], Winogrande [114], TruthfulQA [78], CommonsenseQA [127], PIQA [15], SocialIQA [117], and BoolQ [26]. These benchmarks amount to a total of 41,297 samples. We evaluate all prompt classifiers (from all datasets mentioned above) on the general set after tuning the parameters of the prompt classifiers. For all datasets, we use the test set if the labels are publicly available; otherwise, we use the validation set. For TruthfulQA, we use the MC1 subset for evaluation.
@@ C.3.2 Training a prompt classifier
For all prompt classifiers used for prompt content detection, we choose either RoBERTa-base [82] or Llama-3.1-1B-Instruct [34] as the base model for fine-tuning. The hyperparameters are selected following prior work that improves stability during training [97]. Since in most cases the number of positive samples (the forget samples) is much smaller than the number of negative samples, we reweight the class-wise losses using the inverse frequency. Once the optimal number of epochs is determined, we fine-tune the model again on the combined training and validation sets and use it as the final prompt classifier for inference-time unlearning. We also evaluate all prompt classifiers on the general set outlined in Table 9.
@@
We report the performance of the original prompt classifier, the classifier with simple thresholding, and the classifier with conformal prediction in Tables 5, 6 and 7. We demonstrate that in all settings considered, our best prompt classifiers achieve satisfying performance, reflected by the low false negative rate on the forget set and the low false positive rate on the retain set. On the general set (i.e., the suite of NLP benchmarks), most prompt classifiers have zero false positive predictions, suggesting that the performance of the main LLM on samples irrelevant to the forget set is unlikely to be affected.
C.3.3 Robustness Against Out-of-Distribution Prompts
Perturbation type False Positive Rate (%) False Negative Rate (%)
None 0.0 0.0
Rephrased 0.14 1.5
Adversarial 0.2 1.5
With irrelevant context 0.17 0.5
With jailbreak-like prefix/suffix 1.65 7.52
Keywords and short phrases 0.28 2.51
Table 10: Original TOFU classifier’s false positive and false negative rates for different types of o.o.d. prompts.
Perturbation type False Positive Rate (%) False Negative Rate (%)
None 0.0 0.0
Rephrased 0.03 0.5
Adversarial 0.06 0.75
With irrelevant context 0.0 0.25
With jailbreak-like prefix/suffix 0.08 1.75
Keywords/short phrases 0.0 0.5
(a) TOFU
Perturbation type False Positive Rate (%) False Negative Rate (%)
None 0.00 0.00
Rephrased 0.01 1.00
Adversarial 0.1 1.1
With relevant context 0.00 0.83
With jailbreak-like prefix/suffix 0.07 3.50
Keywords/short phrases 0.00 1.03
(b) WMDP
Perturbation type False Positive Rate (%) False Negative Rate (%)
None 0.0 0.0
Rephrased 0.01 0.2
Adversarial 0.03 1.81
With relevant context 0.00 0.37
With jailbreak-like prefix/suffix 0.09 2.63
Keywords/short phrases 0.0 0.18
(c) Copyrighted content (HP book)
Table 11: Prompt classifiers’ false positive and false negative rates for different types of o.o.d. prompts after being trained on synthetically generated o.o.d. data. The evaluation is still conducted on human-generated o.o.d. data.
In the main body of the paper, we only considered in-distribution prompts from users. However, this assumption is not entirely realistic in practice, as user behaviors are dynamic, and potential attackers could exploit flaws in prompt classifiers to bypass guardrails. Given that the default prompt classifiers we use assume in-distribution prompts and are not trained to identify out-of-distribution (o.o.d.) or jailbreak prompts, we first 1) study the fragility of these classifiers against various types of o.o.d. prompts, and then 2) assess whether the classifiers’ performance improves when trained on these types of prompts.
Original classifiers remain robust under distribution shift.
We consider challenging queries written by humans, including rephrased prompts, adversarial prompts, prompts with irrelevant context, jailbreak prefixes/suffixes, and keyword-only prompts (based on the original prompts). To evaluate, we construct a test set of such prompts by rewriting the prompts from the forget set. Surprisingly, when we evaluate the original TOFU classifier on this test set, it still maintains high false positive and false negative rates, as shown in table 10. This confirms that the classifier’s performance, as reported in table 6, is not due to overfitting. Based on these results, we believe that our classifiers remain robust against common perturbations, even without explicit training on such data.
Training on additional o.o.d. data further improves robustness.
Next, we investigate whether we can enhance the robustness of our classifiers. We construct another set of perturbed prompts, synthetically generated by Llama-3.1-70B-Instruct [3], with no overlap with the previous perturbed test set and no prior knowledge about the perturbation types. Specifically, we prompt Llama-3.1-70B-Instruct to generate a more challenging version of the prompt that could potentially deceive a well-performing classifier. We use this set of prompts to train our classifier and observe that both the false positive rate and false negative rate are further improved (table 11). We repeat this procedure for all classifiers to ensure our findings generalize.